OpenStudy (dominirican1013):

Please help! what is the angle between v=(-7,8) and the positive x-axis? I need to use a formula (which I will write below) to find the cosine of the angle between the two vectors, (-7,8) and (1,0).

OpenStudy (anthonyym):

You want to draw the picture out.

OpenStudy (dominirican1013):

Yes, please.

OpenStudy (anthonyym):

 Do you understand so far?

OpenStudy (anonymous):

\(\textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (dominirican1013):

It's blank I can't see anything